OpenStudy (anonymous):

f(x) = 3 x^(2)ln x what is f'(x)= andf'(e^4)=

OpenStudy (campbell_st):

find the 1st derivative using the product rule u 3x^2 u' = 6x v = ln(x) v' = 1/x so \[f'(x) = u \times v' + v \times u'\] substitute and simplify... the 2nd part is the substitute \[x = e^4\] into the derivative above. hope this helps

OpenStudy (anonymous):

f'(x)= \[3x(1+ 2\ln x )\] f'(e^4)= \[27e ^{4}\]

OpenStudy (anonymous):

thanks :)